OpenStudy (anonymous):

As the particle passes through the point (4,2), its x-coordinate increases at a rate of 7 cm/s. How fast is the distance from the particle to the origin changing at this instant?

OpenStudy (michele_laino):

hint: we have the subsequent position function: \[x\left( t \right) = 7t + C\]

OpenStudy (michele_laino):

where C is a constant to be determined using the initial conditions

OpenStudy (michele_laino):

we know that at t=0, x=4, so we can write: \[\begin{gathered} 4 = x\left( 0 \right) = 7 \cdot 0 + C = C \hfill \\ C = 4 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

so our position fuction will be this: \[x\left( t \right) = 7t + 4\]